OpenStudy (anonymous):

How can you write the expression with a rationalized denominator? Cube root 2 / cube root 6

OpenStudy (anonymous):

where r u getting stuck? or what are you not understanding?

OpenStudy (anonymous):

I'm not understanding any of it.

OpenStudy (anonymous):

suppose: we have: \[\sqrt{3}/\sqrt{5}\] and i want to rationalising it... i will multiply and divide whole expression by such number that denominator becomes a whole number... so here that number is

OpenStudy (anonymous):

\[\sqrt{5}\] so we get ...\[(\sqrt{3}*\sqrt{5})/(\sqrt{5}*\sqrt{5})\]

OpenStudy (anonymous):

now what our expression turns out to be?

OpenStudy (anonymous):

Idk, that's not my problem so therefore I don't see any use in doing it. I wanna know how to do the problem I posted

OpenStudy (anonymous):

if u understand this then you will be able to do ur problem...

OpenStudy (anonymous):

I just wanna do my problem, don't have anymore like this.